# MMMU-Pro: A More Robust Multi-discipline Multimodal Understanding Benchmark

### v2

@@ Abstract
This paper introduces MMMU-Pro, a robust version of the Massive Multi-discipline Multimodal Understanding and Reasoning (MMMU) benchmark. MMMU-Pro rigorously assesses multimodal models’ true understanding and reasoning capabilities through a three-step process based on MMMU: (1) filtering out questions answerable by text-only models, (2) augmenting candidate options, and (3) introducing a vision-only input setting where questions are embedded within images. This setting challenges AI to truly “see” and “read” simultaneously, testing a fundamental human cognitive skill of seamlessly integrating visual and textual information. Results show that model performance is substantially lower on MMMU-Pro than on MMMU, ranging from 16.8% to 26.9% across models. We explore the impact of OCR prompts and Chain of Thought (CoT) reasoning, finding that OCR prompts have minimal effect while CoT generally improves performance. MMMU-Pro provides a more rigorous evaluation tool, closely mimicking real-world scenarios and offering valuable directions for future research in multimodal AI.
Figure 1: Overview of different multimodal LLMs’ performance on MMMU-Pro and MMMU (Val).
@@ 1 Introduction
Recent advances in multimodal large language models (MLLMs) have led to remarkable progress in tackling complex reasoning tasks that combine textual and visual information (Yin et al., 2023a; Jin et al., 2024). Models like GPT-4o (OpenAI, 2024a) have achieved impressive results, e.g., on the Massive Multi-discipline Multimodal Understanding and Reasoning (MMMU) benchmark (Yue et al., 2024), reaching an accuracy of 69.1% on college-level questions that integrate text and images.
@@
This question has profound implications for the development and deployment of AI systems in real-world applications. If models rely on superficial cues rather than true multimodal understanding (Du et al., 2023; Yuksekgonul et al., 2023), we risk overestimating their capabilities and potentially deploying systems that fail in unpredictable ways when faced with novel scenarios (Wu and Xie, 2024; Tong et al., 2024b).
To address this concern and push the boundaries of multimodal AI evaluation, we introduce MMMU-Pro, a more robust and challenging version of the MMMU benchmark. MMMU-Pro is designed to more accurately and rigorously assess a model’s true multimodal understanding and reasoning capabilities across a wide range of academic disciplines. The development of MMMU-Pro is motivated by several key observations, including the text-only solvability of some existing benchmark questions, the limited option space in multiple-choice formats (Wang et al., 2024), and the need to challenge models’ ability to jointly understand different modalities in a more integrated way.
@@
The introduction of the vision-only input setting is particularly crucial, as it tests a fundamental human cognitive ability: the seamless integration and switching between visual and textual information. This setting challenges models to develop the capability to truly “see” and “read” simultaneously, mirroring how humans effortlessly process complex scenes where text and images are intertwined. This ability is crucial for tasks ranging from interpreting scientific diagrams (Li et al., 2024d) to navigating graphical user interfaces (Liu et al., 2024b; Zheng et al., 2024; Koh et al., 2024). Moreover, this approach aligns with how users naturally interact with AI systems, often sharing screenshots or photos rather than meticulously separating text and images.
Our experimental results demonstrate the effectiveness of MMMU-Pro in providing a more rigorous evaluation of multimodal models. We observe significant performance drops across all tested models when compared to the original MMMU benchmark, with decreases ranging from 16.8% to 26.9%. These results highlight the limitations of current state-of-the-art models in true multimodal understanding and reasoning. Furthermore, our analysis reveals that while Chain of Thought (CoT) (Wei et al., 2022) prompting generally improves performance, the benefits vary across models and settings.
@@
Figure 3: Sample questions from MMMU-Pro. In the Vision Input setting, the model is required to answer a multiple-choice question with up to 10 options, each embedded within a screenshot or photo. These images were manually captured by human annotators in diverse display environments to reflect real-world scenarios.
@@ 2.1 Revisiting the MMMU Benchmark
The Massive Multi-discipline Multimodal Understanding and Reasoning (MMMU) benchmark (Yue et al., 2024) is a comprehensive dataset designed to evaluate multimodal AI models on college-level tasks requiring subject-specific knowledge and deliberate reasoning. MMMU consists of 11.5K carefully curated multimodal questions from college exams, quizzes, and textbooks, covering six core disciplines across 30 subjects and 183 subfields. Each question in MMMU is a multimodal image-text pair with 4 multiple-choice options, featuring 30 diverse image types such as charts, diagrams, maps, and chemical structures. MMMU has quickly become a standard evaluation tool in the field, used to assess the capabilities of many prominent multimodal models upon their release (OpenAI, 2024a, b; Anthropic, 2024; Reid et al., 2024; Li et al., 2024b).
Meanwhile, we received some feedback from the community that text-only LLMs could accurately answer some questions without any need for visual input. We took a close look at these questions and identified two main issues: 1) Text-Only Dependency: Certain questions are relatively independent or irrelevant to the corresponding images. 2) Shortcut Exploitation: Even when questions require images for humans to answer correctly, models could often find shortcuts or correlations within the candidate options, leveraging their pre-existing knowledge (from pre-training) to arrive at the correct answer. Two examples that are answered correctly by Llama-3-70B Instruct (Dubey et al., 2024) are shown in Figure 4.
@@ 2.2 Methods
To mitigate the issues and construct a more robust benchmark, we implemented a three-step approach:
Filtering Questions:
We began by filtering out questions that could be answered by text-only LLMs. We selected four strong open-source LLMs: Llama3-70B-Instruct (Dubey et al., 2024), Qwen2-72B-Instruct (Yang et al., 2024), Yi-1.5-34B-Chat (Young et al., 2024), and Mixtral-8$\times$22B-Instruct (gpt-4o, )—and tasked them with answering the MMMU questions without access to images. The models were required to provide answers even when they indicated that visual input was necessary. We repeated this process ten times for each model, considering a question as "answerable" if a model correctly answered it more than five times. We then excluded any question where at least three out of the four models answered correctly across the majority of trials. We then randomly sampled 1800 questions from the remaining pool, evenly distributed across 30 subjects (60 questions per subject).
@@
Augmenting Candidate Options:
Despite the filtering, some questions could still be answered by text-only LLMs, often by exploiting subtle hints within the candidate options. To counteract this, we increased the number of candidate options from four to ten, making it more challenging for models to rely on guessing. This augmentation was done by human experts with the assistance of GPT-4o. During this process, experts also reviewed the original annotated questions to ensure their relevance to the images and to eliminate any questions that lacked a clear connection or coherence. This step filtered out 70 questions and we obtained 1730 questions in total.
@@
As illustrated in Figure 5, these two steps significantly reduced the accuracy of text-only models attempting to guess the answers.
Enhancing Evaluation with a Vision-Only Input Setting:
@@
To further challenge the multimodal understanding of models, we introduced a vision-only input setting in MMMU-Pro. In this setting, the model is presented with a question embedded within a screenshot or photo, without any text explicitly fed into the model. To implement this setting, we asked the human annotators to manually capture photos and screenshots over a simulated display environment. This process involved varying the backgrounds, font styles, and font sizes to replicate the diversity of real-world conditions. By using different combinations of these elements, we created a broad range of visual contexts, ensuring that the models are not only challenged by the integration of text and images but also by the variability in how this content is presented. Examples of the vision-only input setting are shown in Figure 3.
@@
The motivation for introducing this setting stems from real-world usage patterns and the fundamental cognitive abilities of humans. In everyday scenarios, users often take shortcuts by simply capturing screenshots of questions that include both text and images, rather than inputting text separately. This habit of feeding integrated visual-textual content reflects a natural human tendency to process information holistically. Humans excel at interpreting and understanding information when both text and images are presented together, and our goal with this setting is to push models toward achieving a similar level of comprehension. By mimicking this real-world behavior, the vision-only input setting not only adds realism to the benchmark but also ensures that models are better equipped to handle the complexities of multimodal tasks as they would appear in practical applications.
After these three steps, we finally obtained 3460 questions in total (1730 samples are in the standard format and the other 1730 are in the screenshot or photo form).
Table 1: Overall results of different models on different subsets. $\Delta_{1}$: Standard (10 options) - Standard (4 options); $\Delta_{2}$: Vision - Standard (10 options); $\Delta_{3}$: MMMU-Pro - MMMU (Val). MMMU-Pro overall score refers to the average of Standard (10 Options) and Vision Input settings.
MMMU-Pro MMMU (Val) $\Delta_{1}$ $\Delta_{2}$ $\Delta_{3}$
Model
@@
Vision Overall
Random Choice 24.9 12.8 12.4 12.6 22.1 -12.1 -0.4 -9.8
Frequent Choice 27.8 12.1 12.1 12.1 26.8 -15.7 0 -14.7
GPT-4o (0513) (OpenAI, 2024a) 64.7 54.0 49.7 51.9 69.1 -10.7 -4.3 -17.2
Claude 3.5 Sonnet (Anthropic, 2024) 63.7 55.0 48.0 51.5 68.3 -8.7 -6.9 -16.8
Gemini 1.5 Pro (0801) (Reid et al., 2024) 60.6 49.4 44.4 46.9 65.8 -11.2 -5.0 -18.9
Qwen2-VL-72B  (Qwen, 2024) 59.3 49.2 43.3 46.2 64.5 -10.1 -5.9 -18.3
Gemini 1.5 Pro (0523) (Reid et al., 2024) 57.6 46.5 40.5 43.5 62.2 -11.1 -6.0 -18.7
GPT-4o mini  (OpenAI, 2024b) 55.3 39.9 35.2 37.6 59.4 -15.4 -4.7 -21.9
InternVL2-Llama3-76B  (Chen et al., 2024) 55.0 41.9 38.0 40.0 58.3 -13.1 -3.9 -18.4
InternVL2-40B  (Chen et al., 2024) 47.4 36.3 32.1 34.2 55.2 -11.1 -4.2 -21.0
LLaVA-OneVision-72B  (Li et al., 2024b) 52.3 38.0 24.0 31.0 56.8 -14.3 -14.0 -25.8
Qwen2-VL-7B  (Qwen, 2024) 46.6 34.1 27.0 30.5 54.1 -12.5 -7.1 -23.6
InternVL2-8B  (Chen et al., 2024) 42.6 32.5 25.4 29.0 51.2 -10.1 -7.1 -22.3
MiniCPM-V2.6  (Yao et al., 2024) 40.6 30.2 24.2 27.2 49.8 -10.4 -6.0 -22.6
VILA-1.5-40B  (Lin et al., 2024) 46.8 35.9 14.1 25.0 51.9 -10.9 -21.8 -26.9
LLaVA-NEXT-72B  (Liu et al., 2024a) 43.0 31.0 19.2 25.1 49.9 -12.0 -11.8 -24.8
LLaVA-OneVision-7B  (Li et al., 2024b) 42.8 29.5 18.7 24.1 48.8 -13.3 -10.8 -24.7
LLaVA-NeXT-34B  (Liu et al., 2024a) 44.5 30.3 17.2 23.8 48.1 -14.2 -13.1 -24.4
Idefics3-8B-Llama3  (Laurençon et al., 2024) 40.8 30.1 15.6 22.9 46.6 -10.7 -14.5 -23.8
Qwen2-VL-2B  (Qwen, 2024) 34.8 25.3 17.2 21.2 41.1 -9.5 -8.2 -19.9
Phi-3.5-Vision  (Abdin et al., 2024) 37.8 26.3 13.1 19.7 43.0 -11.5 -13.2 -23.3
LLaVA-NeXT-7B  (Liu et al., 2024a) 33.7 19.4 14.6 17.0 35.3 -14.3 -4.8 -18.4
LLaVA-NeXT-13B  (Liu et al., 2024a) 33.9 19.8 14.5 17.2 36.2 -14.1 -5.3 -19.1
3 Experiments
3.1 Experimental Setups
Baselines. To establish a comprehensive understanding of MMMU-Pro’s difficulty and to provide reference points for future research, we evaluated a diverse set of state-of-the-art multimodal models as baselines. These models represent a range of training approaches and capabilities in the field of multimodal AI. Our baseline models include:
Proprietary Models: GPT-4o (0513) and GPT-4o mini, Claude 3.5 Sonnet, and Gemini 1.5 Pro (0801 and 0523 versions). These models represent the cutting edge of multimodal AI capabilities.
Open-source models: We evaluated a range of open-source models, including InternVL2 (8B, 40B, and Llama3-76B versions), LLaVA (OneVision-7B, OneVision-72B, and various NeXT versions), VILA-1.5-40B, MiniCPM-V2.6, Phi-3.5-Vision, and Idefics3-8B-Llama3. These models showcase the current state of publicly available multimodal AI systems. We evaluated these models across three different settings: 1) Standard setting without augmented options (usually 4 options); 2) Standard setting with augmented options (usually 10 options); 3)Vision-only input setting.
We evaluate the models with both Direct and CoT prompts (as shown in Appendix A), and report the higher ones in the overall results. We also discuss the influence of the CoT prompt in subsection 3.4.

### v3

@@ Abstract
This paper introduces MMMU-Pro, a robust version of the Massive Multi-discipline Multimodal Understanding and Reasoning (MMMU) benchmark. MMMU-Pro rigorously assesses multimodal models’ true understanding and reasoning capabilities through a three-step process based on MMMU: (1) filtering out questions answerable by text-only models, (2) augmenting candidate options, and (3) introducing a vision-only input setting where questions are embedded within images. This setting challenges AI to truly “see" and “read" simultaneously, testing a core human cognitive skill of seamlessly integrating visual and textual information. Results show that model performance is substantially lower on MMMU-Pro than on MMMU, ranging from 16.8% to 26.9% across models. We explore the impact of OCR prompts and Chain of Thought (CoT) reasoning, finding that OCR prompts have minimal effect while CoT generally improves performance. MMMU-Pro provides a more rigorous evaluation tool, closely mimicking real-world scenarios and offering valuable directions for future multimodal research.
[…]
MMMU-Pro: A More Robust Multi-discipline Multimodal
Understanding Benchmark
Xiang Yuethanks: Equal contributions. Contact: xyue2@andrew.cmu.edu , Tianyu Zheng11footnotemark: 1,   Yuansheng Ni11footnotemark: 1, Yubo Wang, Kai Zhang, Shengbang Tong, Yuxuan Sun, Botao Yu, Ge Zhang, Huan Sun, Yu Su, Wenhu Chen, Graham Neubig
MMMU Team https://mmmu-benchmark.github.io/#leaderboard
@@ 1 Introduction
Recent advances in multimodal large language models (MLLMs) have led to progress in tackling complex reasoning tasks that combine textual and visual information (Yin et al., 2023a; Jin et al., 2024). Models like GPT-4o (OpenAI, 2024b) have achieved impressive results, e.g., on the Massive Multi-discipline Multimodal Understanding and Reasoning (MMMU) benchmark (Yue et al., 2024), reaching an accuracy of 69.1% on college-level questions that integrate text and images.
@@
This question has profound implications for the development and deployment of AI systems in real-world applications. If models rely on superficial cues rather than true multimodal understanding (Du et al., 2023; Yuksekgonul et al., 2023), we risk overestimating their capabilities and potentially deploying systems that fail in unpredictable ways when faced with novel scenarios (Wu and Xie, 2024).
To address this concern and push the boundaries of multimodal AI evaluation, we introduce MMMU-Pro, a more robust and challenging version of the MMMU benchmark. MMMU-Pro is designed to more accurately and rigorously assess a model’s true multimodal understanding and reasoning capabilities across a wide range of academic disciplines. The development of MMMU-Pro is motivated by key observations, including the text-only solvability of some benchmark questions, limited option space in multiple-choice formats (Wang et al., 2024), and the need to challenge models’ ability to jointly understand different modalities in a more integrated way.
@@
The introduction of the vision-only input setting is particularly crucial, as it tests a fundamental human cognitive ability: the seamless integration and switching between visual and textual information. This setting challenges models to develop the capability to truly “see” and “read” simultaneously, mirroring how humans effortlessly process complex scenes where text and images are intertwined. This ability is crucial for tasks ranging from interpreting scientific diagrams (Li et al., 2024d) to navigating graphical user interfaces (Liu et al., 2024b; Zheng et al., 2024; Koh et al., 2024). Moreover, this approach aligns with how users naturally interact with AI systems, often sharing screenshots or photos rather than separating text and images.
Our experimental results demonstrate the effectiveness of MMMU-Pro in providing a more rigorous evaluation of multimodal models. We observe significant performance drops across all tested models when compared to the original MMMU benchmark, with decreases ranging from 16.8% to 26.9%. These results highlight the limitations of current state-of-the-art models in true multimodal understanding and reasoning. Furthermore, our analysis reveals that while CoT (Wei et al., 2022) prompting generally improves performance, the benefits vary across models and settings.
@@ 2.1 Revisiting the MMMU Benchmark
The Massive Multi-discipline Multimodal Understanding and Reasoning (MMMU) benchmark (Yue et al., 2024) is a comprehensive dataset designed to evaluate multimodal AI models on college-level tasks that require subject-specific knowledge and deliberate reasoning. MMMU consists of 11.5K carefully curated multimodal questions from college exams, quizzes, and textbooks, covering six core disciplines across 30 subjects and 183 subfields. Each question in MMMU is a multimodal image-text pair with 4 multiple-choice options, featuring 30 diverse image types such as charts, diagrams, maps, and chemical structures. MMMU has rapidly established itself as a standard evaluation framework for testing prominent multimodal models upon their release. (OpenAI, 2024b, a; Anthropic, 2024; Reid et al., 2024; Li et al., 2024a).
However, we find that text-only LLMs can accurately answer some questions without requiring any visual input. We take a closer look at these questions and identify two main issues: 1) Text-Only Dependency: Certain questions are relatively independent or irrelevant to the corresponding images. 2) Shortcut Exploitation: Even when questions require images for humans to answer correctly, models often find shortcuts or correlations within the candidate options, leveraging their pre-existing knowledge (from pre-training) to arrive at the correct answer. Two examples that are answered correctly by Llama-3-70B Instruct (Dubey et al., 2024) are shown in Figure 2.
@@ 2.2 Methods
To address these issues and build a more robust benchmark, we implemented a three-step approach.
Filtering Questions: We begin by filtering out questions that can be answered by text-only LLMs. We select four strong open-source LLMs: Llama3-70B-Instruct (Dubey et al., 2024), Qwen2-72B-Instruct (Yang et al., 2024), Yi-1.5-34B-Chat (Young et al., 2024), and Mixtral-8$\times$22B-Instruct (gpt-4o, )—and task them with answering the MMMU questions without access to images. The models are required to provide answers even when they indicate that visual input is necessary. We repeat this process ten times for each model, considering a question as “answerable” if a model correctly answers it more than five times. We then exclude any question where at least three out of the four models answer correctly across the majority of trials. We randomly sample 1800 questions from the remaining pool, evenly distributed across 30 subjects (60 questions per subject).
Augmenting Candidate Options: Despite the filtering, some questions can still be answered by text-only LLMs, often exploiting subtle hints within the candidate options. To counteract this, we increase the number of candidate options from four to ten, making it more challenging for models to rely on guessing. This augmentation is done by human experts with the help of GPT-4o, with additional validation steps to ensure the quality and diversity of the options. Specifically, GPT-4o generates and Claude 3.5 filters the options, followed by two rounds of human review to refine and verify the augmented options. This augmentation is done by human experts with the help of GPT-4o. During this process, experts also review the original annotated questions to ensure their relevance to the images and to eliminate any questions that lack a clear connection or coherence. This step filters out 70 questions, and we obtain 1730 questions in total.
@@
As illustrated in Figure 3, these two steps significantly reduce the accuracy of text-only models attempting to guess the answers.
Figure 4: Sample questions from MMMU-Pro Vision. The model is required to answer a multiple-choice question with up to 10 options, each embedded within a screenshot or photo. The images were manually captured by annotators in diverse display environments to reflect real-world cases.
@@
Enhancing Evaluation with a Vision-Only Setting: To further challenge the multimodal understanding of models, we introduce a vision-only input setting in MMMU-Pro. In this setting, the model is presented with a question embedded within a screenshot or photo, without any text explicitly fed into the model. To implement this setting, we ask the human annotators to manually capture photos and screenshots over a simulated display environment. This process involves varying the backgrounds, font styles, and font sizes to replicate the diversity of real-world conditions. By using different combinations of these elements, we create a broad range of visual contexts, ensuring that the models are not only challenged by the integration of text and images but also by the variability in how this content is presented. Examples of the vision-only input setting are shown in Figure 4.
@@
The motivation for this setting comes from real-world usage and human cognition. Users often capture screenshots of questions with both text and images instead of inputting text separately, reflecting a natural tendency to process information holistically. Humans excel at understanding integrated visual-textual content, and this setting encourages models to develop similar comprehension. By mimicking this behavior, the vision-only input setting enhances realism and prepares models for real-world multimodal tasks. Ultimately, we obtain 3,460 questions—1,730 in standard format and 1,730 as screenshots or photos.
3 Experiments
3.1 Experimental Setups
Baselines. To establish a comprehensive understanding of MMMU-Pro’s difficulty and to provide reference points for future research, we evaluate a diverse set of state-of-the-art multimodal models as baselines. These models represent a range of training approaches and capabilities in the field of multimodal AI. Our baseline models include:
Proprietary Models: GPT-4o (0513) (OpenAI, 2024b) and GPT-4o mini (OpenAI, 2024a), Claude 3.5 Sonnet (Anthropic, 2024), and Gemini 1.5 Pro (0801 and 0523 versions) (Team et al., 2023; Reid et al., 2024). These models represent the cutting edge of multimodal AI capabilities.
Open-source models: We evaluate a range of open-source models, including InternVL2 (8B, 40B, and Llama3-76B versions) (Chen et al., 2024), LLaVA (OneVision-7B, OneVision-72B, and various NeXT versions) (Li et al., 2024a; Liu et al., 2024a), VILA-1.5-40B (Lin et al., 2024), MiniCPM-V2.6 (Yao et al., 2024), Phi-3.5-Vision (Abdin et al., 2024), and Idefics3-8B-Llama3 (Laurençon et al., 2024). These models showcase the current state of publicly available multimodal AI systems. We evaluate these models across three different settings: 1) Standard setting without augmented options (usually 4 options); 2) Standard setting with augmented options (usually 10 options); 3)Vision-only input setting.
We evaluate the models with both Direct and CoT prompts (as shown in Appendix A), and report the higher ones in the overall results. We also discuss the influence of the CoT prompt in 3.3.
MMMU-Pro MMMU (Val) $\Delta_{1}$ $\Delta_{2}$
@@ 3.1 Experimental Setups
Vision
Random Choice 24.9 12.8 12.4 22.1 -9.3 -9.7
Frequent Choice 27.8 12.1 12.1 26.8 -14.7 -14.7
Human Expert (Low) 75.4 73.0 73.0 76.2 -3.2 -3.2
Human Expert (Medium) 82.1 80.8 80.8 82.6 -1.8 -1.8
Human Expert (High) 88.6 85.4 85.4 88.6 -3.2 -3.2
[…]
Gemini 1.5 Pro (0801) (Reid et al., 2024) 60.6 49.4 44.4 65.8 -16.4 ( - ) -21.4 ( - )
Gemini 1.5 Pro (0523) (Reid et al., 2024) 57.6 46.5 40.5 62.2 -15.7 ( - ) -21.7 ( - )
GPT-4o mini (OpenAI, 2024a) 55.3 39.9 35.2 59.4 -19.5 ($\uparrow$ 1) -24.2 ($\uparrow$ 1)
Qwen2-VL-72B (Qwen, 2024) 59.3 49.2 43.3 64.5 -15.3 ( - ) -21.2 ( - )
[…]
InternVL2-40B (Chen et al., 2024) 47.4 36.3 32.1 55.2 -18.9 ( - ) -23.1 ($\downarrow$ 1)
LLaVA-OneVision-72B (Li et al., 2024a) 52.3 38.0 24.0 56.8 -18.8 ( - ) -32.8 ($\uparrow$ 5)
Qwen2-VL-7B (Qwen, 2024) 46.6 34.1 27.0 54.1 -20.0 ($\uparrow$ 1) -27.1 ($\downarrow$ 1)
Pixtral-12B (Mistral, 2024) 47.5 33.4 25.0 52.5 -19.1 ($\uparrow$ 1) -27.5 ( - )
InternVL2-8B (Chen et al., 2024) 42.6 32.5 25.4 51.2 -18.7 ( - ) -25.8 ($\downarrow$ 3)
MiniCPM-V2.6 (Yao et al., 2024) 40.6 30.2 24.2 49.8 -19.6 ($\uparrow$ 1) -25.6 ($\downarrow$ 3)
VILA-1.5-40B (Lin et al., 2024) 46.8 35.9 14.1 51.9 -16.0 ($\downarrow$ 2) -37.8 ($\uparrow$ 9)
LLaVA-NEXT-72B (Liu et al., 2024a) 43.0 31.0 19.2 49.9 -18.9 ( - ) -30.7 ( - )
LLaVA-OneVision-7B (Li et al., 2024a) 42.8 29.5 18.7 48.8 -19.3 ($\uparrow$ 2) -30.1 ($\downarrow$ 1)
LLaVA-NeXT-34B (Liu et al., 2024a) 44.5 30.3 17.2 48.1 -17.8 ($\downarrow$ 2) -30.9 ($\downarrow$ 1)
Idefics3-8B-Llama3 (Laurençon et al., 2024) 40.8 30.1 15.6 46.6 -16.5 ($\downarrow$ 1) -31.0 ( - )
[…]
Phi-3.5-Vision (Abdin et al., 2024) 37.8 26.3 13.1 43.0 -16.7 ( - ) -29.9 ($\uparrow$ 3)
[…]
LLaVA-NeXT-13B (Liu et al., 2024a) 33.9 19.8 14.5 36.2 -16.4 ( - ) -21.7 ($\downarrow$ 1)
[…]
Approximating Human Expert Performance. While rigorous human evaluation of MMMU-Pro provides valuable insights, conducting such an assessment is both time-consuming and costly. Instead, we develop an approach to approximate human expert performance based on the original MMMU human evaluation data. This approximation is justified by several key factors. Firstly, the core content and difficulty of the questions remain unchanged in MMMU-Pro, supporting the validity of using the original human performance data as a close approximation. Secondly, in the original MMMU evaluation, human experts are required to write out their problem-solving processes, significantly reducing the likelihood of random guessing. For questions without detailed solving processes, we randomly select one option from the augmented candidates and recalculate the accuracy. Finally, human experts, with their innate ability to seamlessly integrate visual and textual information, are expected to perform similarly in the vision-only input setting as they do in the original format. Based on these considerations, we posit that human expert performance on MMMU-Pro closely aligns with the original MMMU results, allowing us to maintain a human performance benchmark without incurring the substantial costs of a new expert evaluation. More details of the human estimation performance can be found in Appendix B.